# DS communications: Trends, Challenges, and Solutions

### v3

@@ VI-C Quantum Communications
Some of the quantum properties have gained attention for being used in space [197], DS [198], and terrestrial [197] communications.
@@
The quantum technology is expected to break through the limit of classical technologies in the aspects of security, computational capabilities, as well as information transmissions. Theoretically, quantum cryptography promises to obtain an absolute secure channel. Theoretical and experimental results have shown that QC can increase the capacity of classical channels with methods, for example, superdense coding [199, 200], and superadditivity [201, 200]. Rengaswamy et al. in [202] demonstrated that quantum-optimal joint detection receiver which is based on the belief propagation with quantum messages (BPQM) algorithm [203], through the superadditivity concept, can enhance classical communications. Such a receiver detects complete codewords getting, for BPSK pure-loss channel, to the possibility of attaining Holevo limit, hence becomes suitable for DSOC. In the same direction, [204] showed that joint detection receivers outperform one at a time codeword BPSK demodulators reaching five times classical decoders rate. Kwiat et al. in [200] attempted to make QC possible in DS, taking the advantage of the less lossy channel through superdense coding which enhances the classical communications capacity by a factor of two without the need for preparing entanglement, which is not easy to set up. Realization of superdense teleportation protocol (SDT), proposed in [200] as a candidate in DS and long distance communications in general, which is implementable in a high fidelity, is in progress [205]. Going beyond those accomplishments and establishing a quantum link by entangled photons in DS is still a goal, for example for the Deep Space Quantum Link (DSQL) project [198], which would require high performance QC system to be achieved [206]. Within the next decade, a full quantum teleportation from the Earth to Lunar Gateway, which is a space station that would be orbiting the Moon, though, is expected to be achieved [207].
@@ VI-D Radio Frequency/Free Space Optical Systems
Even though the FSO potentials in DS are tremendous, full implementation of the technology is yet deferred. RF technology is still dominating space communications facilities, as discussed in Section II-C. Some efforts in the literature support the idea of combining RF and FSO in DS as a starting point to emerge the FSO and to get the best of both worlds.
The two schemes for integrating these technologies are hybrid RF/FSO and dual-hop RF/FSO. While a hybrid RF/FSO system combines the RF and FSO technologies in parallel, dual-hop RF/FSO implement them in series. The ability of hybrid RF/FSO to switch between RF and FSO leads to overall DS system performance improvement [208] as it does in terrestrial [209]. In addition, RF/FSO dual-hop method increases coverage, enhances DS communications reliability, and lowers the overall system cost. In [210], for example, the authors demonstrated the benefits of using dual-hub RF/FSO in lowering the probability of link outages in the presence of Solar scintillation. Similarly, authors of [211] analyzed the performance of RF/FSO systems and showed that they outperform both FSO and RF links in weak to moderate coronal turbulence. They also showed that BPSK modulation outperforms some other commonly used schemes in such an environment. [212] has also shown that combining RF/FSO with OAM increases spectral efficiency. To boost capacity and reduce turbulence effects in DS, furthermore, [213] suggested combining OAM multiplexing in a MIMO system, which operates with modified PPM modulation with spatial mode multiplexing in a hybrid RF/FSO system.
There are, besides, dedicated efforts to make RF/FSO operational for DS, the Integrated Radio and Optical Communications (iROC) project [214], for example, features a hybrid RF/FSO aperture antenna, called Teletanna, that reduces SWaP burden and utilizes a beacon-less and high data rate system. iROC is anticipated to be operational around Mars [215] with SWaP comparable to Mars Reconnaissance Orbiter (MRO), and data rate of 267 Mbps and 10 Gbps if the same technology is used for lunar distances.
@@ VI-E Discussions and Outlook
The term IPN is now used to categorize networks that connect a number of spacecrafts placed in different regions of space. For example, a network consisting of a number of Mars and Earth satellites is referred to as a third IPN [216]. However, the envisioned IPN is a network that connects all planets elements analog to the Internet on Earth reliably. The IPN will need to handle the growth of the multimedia data volume [217] and support the human presence in DS.
OAM is getting a lot of attention outside the scope of DS, as being a strong candidate for the upcoming big data era [218]. OAM has a remarkable performance in fiber optics as it can offer a large capacity for long-distance transmission [219]; the number of modes the fiber could handle, for example, can reach 394 [220]. When it comes to wireless communications, nonetheless, OAM performance can be significantly impacted by signal transmission distance, AT, and the ability to detect partially received beams. Hence, although some aspects indicate the interest in this technology in DS [181], more investigation and maturity for the technology will be needed until it becomes applicable to DS.
Among the other advantages QC offer, the improvements they promise in terms of capacity, making them attractive to be implemented in DS to overcome the ultra-long distances losses and meet the demand for the high rate. It will be even of more interest to integrate, for channel multiplexing, OAM with QC, a similar idea was implemented in [221] for a fiber optic link, boosting the capacity even further. Scalable QC techniques are significantly challenging to realize, mainly due to the fact that quantum repeaters, generally speaking, are challenging to deploy [222, 223, 224, 225], as they require high accuracy of quantum measurements, operations, and memory efficiency. Despite the attempts that promise the practicality of quantum repeaters [226] from memory enhancing prospective, interestingly, however, studies have shown deploying QC over Earth-to-space links reaching a distance of 1200 km have been possible without the need of repeaters [227, 228], as they suffer less than their terrestrial counterparts from turbulence. As the turbulence effects are even less between space-to-space links, it would be of interest to expand such studies to the DS.
@@ VII-A Lessons Learned
Although existing RF systems are unable to meet future DS communications demands as opposite to its FSO counterpart, implementing the FSO still faces major impediments. Starting off with the physical layer challenges of DSOC, there are two primary challenges, namely the beam required pointing accuracy and AT. While the former can be tolerated by several techniques for instance, pointing aid tracking beacons, to avoid AT, space-to-space links can be promising. Moreover, modulation, coding, and detection schemes play important rules on mitigating DS environment impairments. PPM modulation schemes is recommended and commonly used schemes for DSOC. Nevertheless, as PPM have performance limitations when high data rates are required, other schemes should be investigated. The detection of a DSOC system significantly affects the capacity limit and modulation performance outcomes. The more sensitive the detector the better it is for DSOC, however sensitivity could inherently be limited by the choice of modulation, coding, and quantum noise. Moving to the networking aspect, DTN is the most recommended architecture in DS mainly defined by the bundle layer and its related addressing mechanism. More attention is needed, though, to enhance the DS protocols to achieve Solar system exploration goals and ultimately universify the network, for which IPN is certainly one of the solutions. Cubsats technology, on the other hand, will be a major contributor for accomplishing the future IPN, we believe. OAM, in addition is a promising technology for achieving data rates in the range of a hundred gigabyte. Not like any other technology, however, there is a gap in the literature in regard to applying OAM in DS. Partial detection, pointing and accuracy, and AT challenges are yet to be overcome. Merging RF and FSO in DS communications, which has been partially applied in DS, could be the starting point for applying FSO. RF/FSO system is shown to preform better than applying FSO alone in a number of cases.
VII-B Conclusion
@@
The success of space exploration missions relies heavily on the reliability and efficiency of DS communications. That is achieved through different technologies and protocols that can overcome the unique challenges of DS. In this paper, we provided an overview of different technologies and protocols and discussed related challenges and suggested solutions. We presented a detailed discussion and outlook regarding covered DSOC physical layer characteristics, physical performance analysis, protocols, and the field advances. We covered the future requirements that need to be fulfilled to cope with the changes in space mission, including compromising between matured existing technologies and recent advances. The paper should provide a basis for researchers interested in learning about and contributing underlying technologies.
@@ Acknowledgement
The authors would like to thank Abderrahmen Trichili, and Bassem Khalfi for their valuable inputs and feedback for the paper.
@@
Binary Phase-Shift Keying
@@
Computer-Generated hologram
@@
Phase-Sensitive Optical Amplifiers

### v4

@@ VI-C Quantum Communications
The quantum technology is expected to break through the limits of classical technologies in the aspects of security, computational capabilities, as well as information transmissions. Quantum technology is marked to provide an absolute secure channel. In addition, theoretical and experimental results have shown that QC can increase the capacity of classical channels with methods such as, superdense coding [210, 211] and superadditivity [212, 211]. Quantum properties have not only gained attention for being used in terrestrial communications [213], but also in space [213] and DS [214]. Interestingly, based on current technology, involving entities from space for achieving terrestrial QC of distances beyond a hundred kilometers is not optional as it provides the means to avoid the hurdles caused by the Earth’s atmosphere. Furthermore, unlike terrestrial mediums where quantum states are susceptible to uncorrected changes [215], because there are negligible channel effects on signals traversing space, it is expected that quantum properties such as entangled photons polarization and other degrees of freedom would persist in DS. The suitability of the DS environment for QC is highlighted in the work [211, 216, 217, 218], where the possibilities of increasing communications rate over DS by the involvement of quantum links are discussed.
@@
Among the ongoing initiatives for establishing DS QC is the Deep Space Quantum Link (DSQL) project [214], which is intended to conduct several unprecedented quantum experiments at long distances including observing gravitational effects on quantum optical systems, and performing the Bell test, teleportation, and entanglement swapping. It is anticipated that within the next decade, a full quantum teleportation from Earth to the Lunar Gateway, which is a space station that would be orbiting the Moon, is expected to be achieved [219]. To realize such objectives, various technological challenges specifically relating to quantum properties that are not yet practical are required to be tackled. The development of DSQL will consequently pave the way for space quantum communications and the realization of a global quantum network [220]. Kwiat et al. [211] attempted to make QC possible in DS, taking advantage of the less lossy channels with superdense coding, which could enhance the classical communications capacity by a factor of two without the need for preparing entanglement. The realization of the superdense teleportation protocol (SDT), proposed in [211] as a candidate for DS, that is implementable in high fidelity is in progress [216]. Rengaswamy et al. [217] demonstrated that a quantum-optimal joint detection receiver, which is based on the belief propagation with quantum messages (BPQM) algorithm [221], through the superadditivity concept can enhance classical communications. Such a receiver detects complete codewords for BPSK pure-loss channel getting to the possibility of attaining the Holevo limit, hence becoming suitable for DSOC. In the same direction, [218] showed that joint detection receivers outperform one at a time codeword BPSK demodulators reaching five times the classical decoders rate.
Furthermore, on a fundamental aspect, there are ongoing studies on the use of the Aharonov-Bohm effect in DS, a quantum phenomenon in which the electromagnetic potential affects particles charge, which promises the possibility of superluminal communications and interplanetary QC. If the application of the effect is realized, the resulting communications capabilities are anticipated to surpass conventional optical communications [222, 223]. Through performing quantum teleportation with X-rays serving as the quantum information carrier, beyond interplanetary links, interstellar communications are possible according to [224, 225]. The aforementioned evidence of the possibility of upgrading DS communications performance beyond the limits of classical methods could render a new set of applications and techniques that mitigate significant DS challenges.
@@ VI-D Radio Frequency/Free Space Optical Systems
Even though the FSO potentials in DS are tremendous, full implementation of the technology is deferred. RF technology is still dominating space communications facilities, as discussed in Section II-C. Some efforts in the literature support the idea of combining RF and FSO in DS as a starting point.
The two schemes for integrating these technologies are hybrid RF/FSO and dual-hop RF/FSO. While a hybrid RF/FSO system combines the RF and FSO technologies in parallel, dual-hop RF/FSO implements them in series. The ability of hybrid RF/FSO to switch between RF and FSO leads to overall DS system performance improvement [226] as it does in terrestrial [227]. The RF/FSO dual-hop method, on the other hand, increases coverage, enhances DS communications reliability, and lowers the overall system cost. In [228], for example, the authors demonstrated the benefits of using dual-hub RF/FSO in lowering the probability of link outages in the presence of Solar scintillation. Similarly, authors of [229] analyzed the performance of RF/FSO systems and showed that they outperform both FSO and RF links in weak to moderate coronal turbulence. They also showed that BPSK modulation outperforms some other commonly used schemes in such an environment. [230] has also shown that combining RF/FSO with OAM increases spectral efficiency. To boost capacity and reduce turbulence effects in DS, furthermore, [231] suggested combining OAM multiplexing in a MIMO system, which operates with modified PPM modulation with spatial mode multiplexing in a hybrid RF/FSO system.
There are, besides, dedicated efforts to make RF/FSO operational for DS. The Integrated Radio and Optical Communications (iROC) project [232], for example, features a hybrid RF/FSO aperture antenna, called Teletanna, that reduces SWaP burden and utilizes a beacon-less and high data rate system. iROC is anticipated to be operational around Mars [233] with SWaP comparable to Mars Reconnaissance Orbiter (MRO) and data rate of 267 Mbps, which is orders of magnitude higher than current rates achievable by the RF technology. Furthermore, if the same technology is used for lunar distances a 10 Gbps data rate can be achieved.
@@ VI-E Discussions and Outlook
The term IPN is now used to categorize networks that connect several spacecrafts placed in different regions of space. For example, a network consisting of several Mars and Earth satellites is referred to as a third IPN [234]. However, the envisioned IPN is a network that connects all planets elements analog to the Internet on Earth reliably. The IPN will need to handle the growth of the multimedia data volume [235] and support the human presence in DS.
OAM is getting a lot of attention outside the scope of DS as a strong candidate for the upcoming big data era. OAM has a remarkable performance in fiber optics as it can offer a large capacity for long-distance transmission [236]; the number of modes the fiber can handle, for example, can reach 394 [237]. When it comes to wireless communications, nonetheless, OAM performance can be significantly impacted by signal transmission distance, atmospheric turbulence, and the ability to detect partially received beams. Resulting possibly in trading off the increase in OAM dimensionality and hence data rate with achieving quality performance. More investigation and maturity for the technology will be needed until it becomes applicable to DS.
Among the other advantages QC offers, the improvements they promise in terms of capacity, making them attractive to be implemented in DS to overcome the ultra-long distance losses and meet the demand for the high rate. It will be even more interesting to integrate, for channel multiplexing, OAM with QC. A similar idea has been implemented for a fiber optic [238] and space [239] links for capacity boosting. Scalable QC techniques are significantly challenging to realize, mainly due to the fact that quantum repeaters, generally speaking, are challenging to deploy [240, 241, 242, 243], as they require high accuracy in quantum measurements and operation and memory efficiency. Despite the attempts that promise the practicality of quantum repeaters [244] and memory enhancement, interestingly, however, studies have shown deploying QC over Earth-to-space links reaching a distance of 1200 km has been possible without the need of repeaters [245, 246], as they suffer less than their terrestrial counterparts from turbulence. As the turbulence effects are even less between space-to-space links, it would be of interest to expand such experimental studies to DS given the fact that utilization of QC over space links is preferred.
@@ VII-A Lessons Learned
Existing RF systems are unable to meet future DS communications demands as opposed to their FSO counterpart. Rendering to the FSO wavelength characteristics, which provide an abundance of bandwidth, low beam divergence and high directivity can be achieved. Consequently, FSO promises significant improvements in system performance including lower SWaP requirements and substantially higher-capacity returns than RF. However, implementing FSO still faces major impediments. Starting off with the physical layer, there are three primary challenges, namely handling the strong power dissipation, pointing loss, and atmospheric effects. While the first can be tolerated by obtaining high-gain transmitters and receivers, pointing loss can be tackled by aid tracking beacons. The space-to-space links can be promising not only to avoid AT effects, but also help to realize OAM, global quantum communications, and IPN.
Modulation, coding, and detection schemes play important roles in mitigating DS environment impairments. As concluded, PPM scheme is highly recommended and commonly used among the DSOC research community. PPM has a high peak-to-average power ratio which leads to capacity improvement. Nevertheless, PPM has performance limitations when higher-data rates are demanded. In addition, increasing PPM order results in higher-system complexity and cost. Furthermore, the capacity limit analysis suggests to investigate other modulation schemes than PPM. When it comes to coding, methods that are well studied in DS are limited and mostly restricted to PPM. The detection of a DSOC system significantly affects the capacity limit and modulation performance outcomes. The more sensitive the detector the better it is. However, sensitivity could inherently be limited by the choice of modulation, coding, and quantum noise. Although the trend recommends direct detection techniques, coherent detection methods are gaining more attention currently as they achieve higher-information rates.
@@
Moving to the networking aspect, DTN is the most recommended architecture in DS mainly defined by the bundle layer and its related addressing mechanism. However, more attention is needed to enhance the DS protocols to achieve Solar system exploration goals and ultimately universify the network, for which, IPN is certainly one of the solutions. On the other hand, we believe that Cubesats technology will be a major contributor to accomplishing the future IPN. Moreover, OAM is a promising technology for achieving data rates around hundreds of gigabytes. Unlike other technologies, there is a gap in the literature in terms of applying OAM in DS. Moreover, partial detection, pointing and displacement, and AT challenges are yet to be overcome. Acknowledging that the objectives to realize a global quantum network could be achieved by utilizing space links should motivate more investment in DS QC, which in turn could provide a new perspective on the technology employment in general. Merging RF and FSO in DS communications, which has been partially applied in DS, could be the starting point for applying FSO. RF/FSO systems are shown to perform better than FSO in a number of cases.
VII-B Summary
@@
The success of space exploration missions relies heavily on the reliability and efficiency of DS communications. That is achieved through different technologies and protocols, which can overcome the unique challenges of DS. In this paper, we provided an overview of different technologies and discussed related challenges and possible solutions. We presented a detailed discussion on the DSOC physical layer characteristics, physical performance analysis, protocols, and field advances. We addressed the future requirements that need to be fulfilled to cope with the changes in space missions, including compromising between mature existing technologies and recent advances. The paper should provide a basis for researchers interested in learning about and contributing to the underlying technologies.
@@ Acknowledgement
The authors would like to thank Abderrahmen Trichili, and Bassem Khalfi for their valuable input and feedback for the paper.
@@
Binary Phase Shift Keying
@@
Computer Generated Hologram
@@
PIE
Photon Information Efficiency
PNR
Photon Number Resolved
@@
Phase Sensitive Optical Amplifiers